# LMV-RPA: Large Model Voting-based Robotic Process Automation

### v1

@@ 5 Results and Discussion
Table 1 summarizes the performance comparison of our custom MLV-RPA model against the state-of-the-art RPA tools, UiPath and Automation Anywhere across different tasks. The evaluation highlights both the average runtime and accuracy for each model.
@@
MLV-RPA (Ours) 121.27 sec
Table 1, we observe a comparison of automation models based on their average run times. UiPath and Automation Anywhere exhibit relatively longer average run times, clocking in at 212.33 seconds and 217.80 seconds, respectively. In contrast, MLV-RPA (Ours) demonstrates a significantly shorter average runtime of 121.27 seconds.
Despite UiPath and Automation Anywhere being well-established platforms with substantial capabilities, MLV-RPA’s shorter runtime showcases its efficiency in processing tasks faster. The reduced runtime of MLV-RPA can be attributed to its optimized architecture, which leverages advanced techniques to maintain high performance while minimizing execution time.
Overall, MLV-RPA stands out as a more efficient solution, offering a significant improvement in processing speed compared to UiPath and Automation Anywhere. This makes it particularly suitable for scenarios where minimizing execution time is critical without sacrificing precision.
@@
Our MLV-RPA (4 OCR models + 2 layers LLM with Majority Voting) 99%
Table 2 presents a comparison of automation models based on accuracy. The traditional model, which uses 4 OCR engines combined with a single-layer LLM, achieves an accuracy of 94%. However, the MLV-RPA model (our proposed approach), which incorporates 4 OCR engines and a two-layer LLM with a majority voting mechanism, achieves a significantly higher accuracy of 99%.
@@ 6 Conclusion
This study has provided a comprehensive comparison of UiPath, Automation Anywhere, MLV-RPA, focusing on how these three tools are able to execute the automation processes quicker and more accurately. While Both UiPath and Automation Anywhere show regular performance and require longer processing time; the custom-built model of MLV-RPA proved to be very accurate with a result of 99%, although taking more time in processing tasks. By contrast, LMV-RPA strikes the best balance of speed and accuracy. Thus, it is an extremely effective solution for quick and reliable automation.
These findings underline the potential of custom-built RPA systems in solving particular automation challenges. Indeed, MLV-RPA is best for activities that demand high accuracy, whereas LMV-RPA works best for those tasks that require the highest speed. Hence, it is vital for organizations to consider their needs for automation and balance speed and accuracy requirements while choosing an RPA tool to improve efficiency in operations and meet their business objectives successfully.

### v2

@@ 5 Results and Discussion
Table 1 summarizes the performance comparison of our custom LMV-RPA model against the state-of-the-art RPA tools, UiPath and Automation Anywhere across different tasks. The evaluation highlights both the average runtime and accuracy for each model.
@@
LMV-RPA (Ours) 121.27 sec
Table 1, we observe a comparison of automation models based on their average run times. UiPath and Automation Anywhere exhibit relatively longer average run times, clocking in at 212.33 seconds and 217.80 seconds, respectively. In contrast, LMV-RPA (Ours) demonstrates a significantly shorter average runtime of 121.27 seconds.
Despite UiPath and Automation Anywhere being well-established platforms with substantial capabilities, LMV-RPA’s shorter runtime showcases its efficiency in processing tasks faster. The reduced runtime of LMV-RPA can be attributed to its optimized architecture, which leverages advanced techniques to maintain high performance while minimizing execution time.
Overall, LMV-RPA stands out as a more efficient solution, offering a significant improvement in processing speed compared to UiPath and Automation Anywhere. This makes it particularly suitable for scenarios where minimizing execution time is critical without sacrificing precision.
@@
Our LMV-RPA (4 OCR models + 2 layers LLM with Majority Voting) 99%
Table 2 presents a comparison of automation models based on accuracy. The traditional model, which uses 4 OCR engines combined with a single-layer LLM, achieves an accuracy of 94%. However, the LMV-RPA model (our proposed approach), which incorporates 4 OCR engines and a two-layer LLM with a majority voting mechanism, achieves a significantly higher accuracy of 99%.
@@ 6 Conclusion
This study has provided a comprehensive comparison of UiPath, Automation Anywhere, LMV-RPA, focusing on how these three tools are able to execute the automation processes quicker and more accurately. While Both UiPath and Automation Anywhere show regular performance and require longer processing time; the custom-built model of LMV-RPA proved to be very accurate with a result of 99%, although taking more time in processing tasks. By contrast, LMV-RPA strikes the best balance of speed and accuracy. Thus, it is an extremely effective solution for quick and reliable automation.
These findings underline the potential of custom-built RPA systems in solving particular automation challenges. Indeed, LMV-RPA is best for activities that demand high accuracy, whereas LMV-RPA works best for those tasks that require the highest speed. Hence, it is vital for organizations to consider their needs for automation and balance speed and accuracy requirements while choosing an RPA tool to improve efficiency in operations and meet their business objectives successfully.